OpenStudy (anonymous):

if f(x) = x − 4 and g(x) = 1/2 x − 2, find f[g(−6)]. 9 3 −9 −3

jimthompson5910 (jim_thompson5910):

are you able to find g(-6) ?

OpenStudy (anonymous):

no i havent been thought thislesson this is th epretest that i was sent for homework

OpenStudy (anonymous):

tought

jimthompson5910 (jim_thompson5910):

g(x) = 1/2 x − 2 g(-6) = (1/2)(-6) − 2 ... replace every x with -6 g(-6) = -6/2 − 2 g(-6) = -3 − 2 g(-6) = ???

OpenStudy (anonymous):

-5?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so g(-6) = -5

jimthompson5910 (jim_thompson5910):

this means f[g(−6)] is the same as f(-5) after you replace the "g(-6)" with -5 (they are the same, so you can do this)

jimthompson5910 (jim_thompson5910):

now what is f(-5) ?

OpenStudy (anonymous):

i have no idea honestly

jimthompson5910 (jim_thompson5910):

we are given f(x) = x − 4

jimthompson5910 (jim_thompson5910):

to find f(-5), we just replace each x with -5 and evaluate f(x) = x − 4 f(-5) = -5 − 4 f(-5) = ???

OpenStudy (anonymous):

-9

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that's your final answer because f(-5) = f[g(−6)]

OpenStudy (anonymous):

i understand it now thank you sooo much!

jimthompson5910 (jim_thompson5910):

you're welcome